OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 3 = -12x to the nearest hundredth?

OpenStudy (anonymous):

i give medals

OpenStudy (anonymous):

here are the options x ≈ -3.23 and x ≈ 0.23 x ≈ -2.72 and x ≈ -0.28 x ≈ 0.28 and x ≈ 2.72 x ≈ -0.23 and x ≈ 3.23

OpenStudy (ranga):

It is a quadratic equation. 4x^2 + 3 = -12x 4x^2 + 12x + 3 = 0 Use the quadratic formula: { -b +/- sqrt(b^2 - 4ac) } / 2a

OpenStudy (anonymous):

so it would be x=2.72 and x=1.40??

OpenStudy (ranga):

I am getting x = -2.72 and x = -0.28

OpenStudy (anonymous):

show me how you got that ??

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

can you show me how to solve for x?

OpenStudy (ranga):

4x^2 + 12x + 3 = 0 x = { -12 +/- sqrt( 12^2 - 4*4*3 ) } / (2*4) x = { -12 +/- sqrt(144-48) } / 8 x = { -12 +/- sqrt(96) } / 8 x = { -12 +/- 9.8 } / 8 x = (-12 + 9.8) / 8 or (-12 - 9.8) / 8 = -2.2/8 or -21.8 / 8 x = -0.28 or -2.72